OpenStudy (anonymous):

A computer company gives a discount of 10% on a new piece of software. If it makes a profit of 4%, what percent does the marked price exceed the cost price?

OpenStudy (anonymous):

p=original price d=discount d=.9p P=profit P=.04d

OpenStudy (anonymous):

it is near to $ 4

OpenStudy (anonymous):

14%

OpenStudy (anonymous):

cost price = .96d

OpenStudy (anonymous):

13.6%

OpenStudy (anonymous):

I'm not sure...its the last question on my hw and I'm sooo confused!

OpenStudy (anonymous):

cost price is 86.4% of the original

OpenStudy (anonymous):

I'm not even sure where to start with this problem... :-/

OpenStudy (anonymous):

so if the discount is 10%, then it is 90% the original cost. or original cost *.9 if they make 4% profit off the discount price. cost price is 96% of the discount price.

OpenStudy (anonymous):

yeah tutor is right in approximate terms i was sayin but in exact it is\ = 0.9 * 0.96 * 100 = 86.4 therefore 100 - 86.4 = 13.6%

OpenStudy (anonymous):

so if d=.9p, c=.96d, c=.96(.9p)=.864p

OpenStudy (anonymous):

does that make more sense?

OpenStudy (anonymous):

honestly...some but not all. math is a bad subject for me, you all are explaining right I'm sure, I'm just slow.

OpenStudy (anonymous):

hey let us suppose the marked price is $ 100

OpenStudy (anonymous):

then after allowing discount the software will be sold at what price ??

OpenStudy (anonymous):

this equation requires you to think of the flip side. if it is a 10% discount, then you are paying 90% of the price. If it is a 4% profit, then the cost is 96%

OpenStudy (anonymous):

reply i m waiting???

OpenStudy (anonymous):

wait tutor let me explain it first then u give another try

OpenStudy (anonymous):

its the 4% part I'm stuck at...

OpenStudy (anonymous):

tell one thing can we assume the price of software as $100

OpenStudy (anonymous):

say i m waiting

OpenStudy (anonymous):

Light bulb!!! I got it! Sorry I have to read things a few times to understand...sorry slow...Thank you!!

OpenStudy (anonymous):

ok will explain it to u let us first assume the Marked Price of Software is $ 100 (we can assume any value ok)

OpenStudy (anonymous):

got it ???

OpenStudy (anonymous):

Yes I do! Thank you!

OpenStudy (anonymous):

then after allowing a discount of 10 % the software will be sold at what price ?

OpenStudy (anonymous):

90%

OpenStudy (anonymous):

no express this in Dollars

OpenStudy (anonymous):

u ahve assumed the marked price as $ 100 so after allowing a discount of 10% it will get reduced to????

OpenStudy (anonymous):

say???

OpenStudy (anonymous):

90% of $100 = ???

OpenStudy (anonymous):

did you got my feelings?

OpenStudy (anonymous):

$90

OpenStudy (anonymous):

If the original price is $100, they would have to mark it up 10% to earn a profit of 4%. 14% above the original cost...right?